You invest $25,000 today at 8% per year. How much money will you have accumulated after...

  1. You invest $25,000 today at 8% per year. How much money will you have accumulated after 13 years?
  2. You are going to receive $150,000 in 25 years. Calculate the present value of the $150,000 using discount rates of 8% and 10%.
  3. Your friend has learned that he is going to receive $7,500 a year for the next 10 years. Utilizing 5% interest rate, calculate the current value of the future payments.
Answer #1

1)

Principal invested = $ 25,000

Rate = 8%

Number of years = 13 years

Amount after 13 years = 25,000*(1+8%)^13 = $ 67,990.59

2)

Future value in 25 years = 150,000

PV at 8% = 150,000/(1.08)^25 = $ 21,902.68

PV at 10% = 150,000/(1.10)^25 = $ 13,844.39

3)

Amount every year , PMT = 7500

Number of years, nper = 10 years

Rate = 5%

Present value of future payments = J =pv(5%,10,7500 PV(rate, nper, pmt, [fv], [type]) = $ 57,913.01